OpenStudy (anonymous):

sophia has 2 purple flowers, 6 pink flowers and 4 yellow flowers. what fraction of sophias flowers are not yellow.

OpenStudy (anonymous):

total flowers = 2+6+4 = 12 flowers that aren't yellow = 2+6 = 8 fraction that aren't yellow = 8/12 or 2/3

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Most welcome